OpenStudy (anonymous):

How do I begin to take the limit of (4x)^(1/x)?

zepdrix (zepdrix):

Limit as x approaches what value?

OpenStudy (anonymous):

Oh infinity. sorry about that.

zepdrix (zepdrix):

So our function is approaching the `indeterminate form`: \(\Large\sf \infty^{0}\)

zepdrix (zepdrix):

So we need to try and do something fancy to see what's going on. It's not in the correct form to apply L'Hopital just yet. Maybe we can fix that though.

zepdrix (zepdrix):

Here is a handy trick we can use:\[\Large\sf x=e^{\ln x}\]

zepdrix (zepdrix):

\[\Large\sf \lim_{x\to\infty}4x^{1/x}\]So we'll apply this idea to our problem, both exponentiate and take the log,\[\huge\sf e^{\ln(\lim_{x\to\infty}4x^{1/x})}\]We can pull the limit operator out of the log,\[\huge\sf e^{\lim_{x\to\infty}\ln(4x^{1/x})}\]Let's ignore the e base for a moment and just pay attention to what's going on up in the exponent.\[\Large\sf \lim_{x\to\infty}\ln(4x^{1/4})\]

zepdrix (zepdrix):

Woops typo on that last line*\[\Large\sf \lim_{x\to\infty}\ln(4x^{1/x})\]

zepdrix (zepdrix):

Rules of exponents allows us to bring the 1/x outside of the log, yes?

zepdrix (zepdrix):

\[\Large\sf \lim_{x\to\infty}\frac{1}{x}\ln(4x)\]

zepdrix (zepdrix):

To be able to apply L'Hopital, we want something of the form:\[\Large\sf \lim_{x\to\infty}\frac{f(x)}{g(x)}=\frac{0}{0},\qquad or\qquad \lim_{x\to\infty}\frac{f(x)}{g(x)}=\frac{\infty}{\infty}\]

zepdrix (zepdrix):

Ok I'll stop talking for a sec :) What you thinking?

OpenStudy (anonymous):

Okay. Yeah there is a bit of fancy footwork going on there. I'm not awfully familiar with infinity to the zero problems so I'm just looking over what you've got so far.

OpenStudy (anonymous):

So you put the whole limit inside of the natural log when you do e to the ln?

zepdrix (zepdrix):

Yah I'm applying this idea,\[\Large\sf \color{orangered}{x}=e^{\ln \color{orangered}{x}}\]To our problem,\[\Large\sf \color{orangered}{\lim_{x\to\infty}4x^{1/x}}=e^{\ln(\color{orangered}{\lim_{x\to\infty}4x^{1/x}})}\]

OpenStudy (anonymous):

Okay I got you. So next you would use L'hopitals rule on \[\frac{ \ln 4x }{ x }\] ?

zepdrix (zepdrix):

Oh the 4 is in brackets with the x? Ah I've been a little sloppy with my notation, sorry about that. Shouldn't change anything though.

zepdrix (zepdrix):

\[\Large\sf \lim_{x\to\infty}\frac{\ln(4x)}{x}\]is giving us the indeterminate form: \(\Large\sf \dfrac{\infty}{\infty}\) Ah yes, good! Looks like we can apply L'Hop from here!

OpenStudy (anonymous):

Yeah so it would end up being \[\frac{ 4 }{ x }\] and then \[\frac{ 0 }{ 1 }\] right?

zepdrix (zepdrix):

\[\Large\sf \lim_{x\to\infty}\frac{\ln(4x)}{x} \qquad L'H\to \qquad \lim_{x\to\infty}\frac{4}{x}\]This limit is approaching 4/infty, which does not allow us to apply L'Hop a second time. This is no longer an indeterminate form.

zepdrix (zepdrix):

Is that what you were trying to do maybe?

OpenStudy (anonymous):

Really it was a sequence that I was trying to find the limit of. So in that case it diverges?

zepdrix (zepdrix):

No. We have a fraction. The numerator is constant, the denominator is getting really really big. So as a whole, the fraction is approaching zero, right?\[\Large\sf \frac{4}{2bajillion}\approx 0\]

zepdrix (zepdrix):

It turns out that by applying L'Hop a second time, you would have simply gotten lucky. But it wasn't correct to do so, so you gotta be careful! :)

OpenStudy (anonymous):

Hahah alright. I usually make that mistake. Thanks a bunch!

zepdrix (zepdrix):

But also recall this only takes care of the exponent part of our problem. We still have the base e, right?

OpenStudy (anonymous):

That's what I was thinking. Apparently the limit is actually 1 when I throw it into wolfram alpha. so e to the 0 is 1?

OpenStudy (anonymous):

Because we took the e out?

zepdrix (zepdrix):

Yes good, so the whole limit mess simplified to zero. We're left with e^0. Which simplifies to 1, yes.

OpenStudy (anonymous):

Awesome Okay thanks for your help!

zepdrix (zepdrix):

np!